OpenStudy (anonymous):

Is this correct? Problem: x(5x-9)=14 My answer: 14, 9/5.

OpenStudy (anonymous):

We would first distribute the x and move the 14 to the other side 5x^2 - 9x - 14 = 0 Now we factor (5x - 14)(x + 1) = 0 and solve x = 14/5 and -1 I think your method did not work because you did not get all of the terms on one side. I think you left 14 on the other side. We need zero over there on that side.

OpenStudy (anonymous):

my bad ^^ thanks for showing me that

OpenStudy (anonymous):

its 14/5 and its aslo can -1

OpenStudy (anonymous):

so you are wrong

OpenStudy (anonymous):

ok